The phase velocity v of wave in a certain
medium is given by v=a+bX where X is
the wavelength and a and b are constants.
What is its group velocity ?
1
Expert's answer
2019-05-23T09:41:51-0400

The phase velocity of wave

"v=a+b\lambda"

The wave number


"k=\frac{2\pi}{\lambda}"

So, the phase velocity

"v=a+\frac{2\pi b}{k}"

From the other side


"v=\frac{\omega}{k}"

Therefore the frequency

"\omega=vk=ak+2\pi b"

The group velocity


"v_g=\frac{d\omega}{dk}=a"